5(x+1)+5(x-1)=26


1
Expert's answer
2020-11-17T06:28:59-0500

This is a typical one-variable equation - "5(x+1)+5(x-1)=26 \to 5x+5+5x-5=26 \to10x=26\to x=\frac{26}{10}\to x=\frac{13}{5}"

Answer: "x=\frac{13}{5}"